Gunpowder contains 75 % nitre and 10 % sulphur .find the amount of gunpowder which carries 9 kg nitre.what amount of gunpowder would contain 2.5 kg sulphur?
1
Expert's answer
2013-07-09T09:58:15-0400
Let x kg be the amount of gunpowder which carries 9 kg nitre. Because gunpowder contains 75% nitre we can say:
x * 0.75 = 9
x = 9 / 0.75
x = 12.
Hence 12 kg gunpowder carries 9 kg nitre.
Let y kg be the amount of gunpowder which contains 2.5 kg sulfur. As gunpowder contains 10% sulfur we obtain:
y * 0.1 = 2.5
y = 25.